### v1

@@ 1 Introduction
We propose MutatiOn TestIng with Fuzzing (MOTIF), an approach that automatically generates fuzz drivers, leverage the data produced by fuzzers, and exercise a mutated function and the corresponding original function, looking for diverging outputs to detect mutants. MOTIF automatically generates seed files for fuzzers and integrates strategies to process mutant-killing inputs to eliminate false positives due to nondeterminism.
Instead of designing a dedicated fuzzing algorithm, MOTIF leverages state-of-the-art (SOTA) fuzzers because we believe that the fuzzers’ bucket-based coverage strategy can effectively select test inputs that eliminate mutants. This strategy not only selects inputs that lead to different software states but also tracks and is guided by the behavioral differences between the original and mutated functions. Although MOTIF can target any C/C++ program, we designed it taking into account characteristics that are prominent in embedded software for CPS. Specifically, CPS software is not executed from the textual console but directly within the operating system, largely relies on floating point instructions, and leverages specific OS and libraries that complicates its analysis or symbolic interpretation. For these reasons, we present MOTIF as a tool for C/C++ embedded software for CPS.
@@
  • Performing a large empirical study to investigate (RQ1) the best fuzzing configurations (compiler, sanitizers, and coverage metric) for mutation testing. Note that the identification of the best fuzzing configuration for different application contexts remains an open problem EURECOM7801 . Our experiments show that the best results are obtained when relying (1) on the Clang compiler, which leads to the quickest fuzz drivers, (2) the address sanitizer (ASAN), which prevents false positives due to violation of function preconditions, and (3) the LAF-Intel LAF:Blog optimization for coverage, which enables computing a better fitness score for mathematical functions, which are, in turn, among the most critical software components. Specifically, the best fuzzing configuration, compared to a standard fuzzing configuration (i.e., AFL++ aflchurn with the GCC GCC compiler), leads to an improvement varying between 1.53 and 4.34 percentage points. Although limited, such improvements may still enable the detection of critical faults; for example, when reusing the generated test cases for regression testing. Further, our findings help engineers make informed decisions on projects setups. For example, mutation testing improvements may not justify a large development endeavor to enable a project to be recompiled with Clang. However, when compiling projects with Clang, ASAN, and LAF-Intel is feasible (e.g., it is sufficient to change a few configuration parameters), the best configuration is a better choice (e.g., enables testing additional boundary cases) and might be considered also for other fuzzing applications beyond mutation testing.
  • Assessing (RQ4) the performance gain achieved by relying on hybrid-fuzzing, which combines fuzzing and symbolic execution. Our results confirm the findings of related work that demonstrate that hybrid fuzzing leads to improved test effectiveness; however, in our experiments, its applicability is limited as it helps significantly improve the mutation score only one in four subjects. The key advantage observed with hybrid-fuzzing is the reduction of killing time (up to 32% faster), which suggests using hybrid-fuzzing for large projects leading to a large number of mutants.
  • Evaluating (RQ5) the cost-effectiveness of reusing mutant-killing inputs to test live mutants; our results suggest that reusing mutant-killing inputs to kill other mutants is recommended when there is a sufficient test budget of at least 5000 seconds per mutant.
@@ 3.4 Step 4 – Perform mutation testing
In Step 4, MOTIF runs the fuzzer to generate inputs for the executable fuzzing driver; in our experiments we leverage the AFL++ fuzzer but the approach is generic and can work with any grey-box fuzzer for C/C++. The fuzzer keeps generating input files until it reports one or more crashes, after which MOTIF halts the fuzzer. This process leads to the generation of fuzzing driver logs and crashing inputs (i.e., input files that caused a crash during the execution of the fuzzing driver). Since fuzzers generate several input files from each input taken from the file pool, and since all of them are executed by the fuzzer, more than one crashing input may be reported.
@@
MOTIF retains all test inputs that kill a mutant. However, the function under test may produce non-deterministic outputs meaning that, despite observed differences in outputs, the inputs may not have actually killed the mutant. For instance, two consecutive invocations of a function that reads and writes global variables may lead to different outputs even if the mutated statement is not exercised; consequently, the input suggested by the fuzzer would be a false positive. To minimize false positives, MOTIF automatically re-executes every test input killing a mutant using a modified version of the fuzzing driver that invokes the original function instead of the mutated function. If this false positive driver, as we refer to it, reports a difference in outputs of the two function calls, it implies that the function under test is non-deterministic and thus that the input may not kill the mutant. MOTIF considers mutants exclusively killed by false positive inputs to be live. To kill mutants in functions that modify global state variables, the user must manually add, in the fuzzing driver between the two function calls, instructions to reset each state variable, similar to other fuzzing approaches for unit and library testing (e.g., LibFuzzer LibFuzzer ).
@@ 3.5 Step 5 – Test the SUT
In this Step, MOTIF generates a test driver for the SUT. An example test driver for function T_POS_IsConstraintValid is shown in Listing 3. The test driver is identical to the fuzzing driver except that (1) it only invokes the original function (Line 5 in Listing 3), (2) instead of comparing outputs from two function invocations, it prints the output data generated by the original function (Lines 7 to 9), and (3) it includes assertions that compare the execution results of the original function with the expected results (expected results are the ones observed during fuzzing for the original function). The test driver is used to test the original SUT with the inputs that kill the mutant and the outputs should then be verified by a software engineer based on the SUT specifications. If the observed output values are correct, they can be used as oracles for future regression testing, which is the reason why the test driver already includes automatically generated test assertions (i.e., MOTIF generates a complete regression test case). Otherwise, a fault has been identified in the SUT. This scenario highlights one of the key advantages of mutation testing: When testing the SUT with inputs that detect simulated human mistakes (mutants), actual faults in the SUT are more likely to be discovered than with randomly selected inputs (e.g., in our experiments we discovered five faults in our subjects).
1int main(int argc, char** argv){
@@ 4 MOTIF C++ extensions
15 return getCharge()/getAverageConsumption();
@@
27 double getAverageConsumption(){
28 return super::getAverageConsumption()-getAverageCharging();
@@ 32 ChargeableModule(double consumptionBoundary, double chargeBoundary){
33 this.consumptionBoundary = consumptionBoundary;
34 this.chargeBoundary = chargeBoundary;
@@ 55 StaticMission(double presetDuration){
64template <class ObjectiveT> class DynamicMission {
@@
Inheritance enables a class to reuse (i.e., inherit) the methods declared by its superclass, if any. Inheritance complicates the unit testing automated by MOTIF-C because it leads to dependencies between superclasses and subclasses. Indeed, not only a subclass may execute superclass methods but the execution of a method declared by a superclass may trigger the execution of a method in a subclass. For example, the execution of method ChargeableModule.isAutonomous() in Listing 4 triggers the execution of the superclass method SatelliteModule.getAutonomy(), which, in turn, triggers the execution of ChargeableModule.getAverageConsumption() and ChargeableModule.getAverageCharging(). As a consequence, the solution adopted by MOTIF-C to test mutants (i.e., creating a copy of the original method and renaming it) would prevent the testing of methods following such pattern; for example, renaming a mutant of method getAverageConsumption() in ChargeableModule as mut_getAverageConsumption() would prevent its call from method SatelliteModule.getAutonomy(), thus rendering testing ineffective.
@@
In MOTIF::TestUtil, methods call and compare take care of the execution of the method under test and the comparison of the results obtained by the original and the mutated method, respectively. Method call (Lines 13-25, Listing 8), which executes the method under test, includes automatically generated instructions to populate parameters with fuzzed data (Lines 17 and 18). If the constructor of a function under test receives as input an abstract type, the instructions needed to construct the concrete type to be used shall be manually added (Lines 22-24). Last, method call invokes the method under test and stores the returned value (Line 26). Method compare, instead, is invoked by the fuzzing driver (Line 13, Listing 7) to compare the results obtained by the original and the mutated class after the execution of the mutant; its logic follows the one of the fuzzing driver generated by MOTIF-C. Approaches for Java MU2 ; fraser2011mutation can leverage programming conventions and reflection to compare object states by leveraging the equals method, which is normally overridden to compare two instances of a class, or inspector methods, which can be identified through naming conventions or purity analysis (simpler to achieve on Java bytecode Salcianu2005 ).
To deal with protected and private methods (i.e., encapsulation), instead of modifying the class under test to specify the test utility class as a friend class, we suggest the end-user to exercise, in the fuzzing driver, a public method that invokes the private/protected method under test. For example, to test mutant 2 in Listing 6, the end-user may still rely on the fuzzing driver shown in Listing 7, because it tests method isAutonomous, which, in turn, executes method getAverageConsumption (i.e., the mutated method). In Java, protected and private methods can be invoked through reflection Arcuri2017 .
@@
5 double constructor_1;
6 double constructor_2;
@@ 13 TypeA object;
17 get_value(&constructor_1, sizeof(constructor_1) );
18 get_value(&constructor_2, sizeof(constructor_2) );
19
20 //instantiate the class under test
21 this->object = new TypeA(param1, param2);
@@
24 this->parameter_1 = new StaticMission(manual_1);
@@
26 this->_return = object->isAutonomous(this->parameter_1);
@@
32 ret += compare_value(this->parameter_1,rhs->parameter_1);
33 ret += compare_value(this->_return,this->_return);
@@ 4.1 Manual tuning of MOTIF’s test utility class
Like MOTIF-C, MOTIF-C++ automatically identifies the variables to create in the test utility class based on method signatures. However, MOTIF-C++ requires the manual specification of the class to instantiate to deal with parametric and subtype polymorphism. For example, if a method declares parameters of an abstract type (e.g., isAutonomous receives one parameter of the abstract type Mission), the end-user needs to specify the additional parameters required to instantiate the concrete class to be used to exercise the method under test (Line 9).
@@ 5.1 Subjects of the study
Our software subjects are provided with test suites whose code coverage is reported in Table 3. Most test suites do not achieve 100% statement coverage because they include components that are tested with specific hardware not available to us; therefore, we generated mutants only for the covered statements. We generated mutants with MASS Oscar:TSE ; MASSTOOL ; specifically, we rely on all the mutation operators supported by MASS (i.e., the sufficient set offutt1996experimental and the deletion set delamaro2014designing ; andrews2005mutation ), which proved effective in previous experiments on similar subjects. We excluded mutants that are identified as equivalent or duplicate according to trivial compiler equivalence methods Oscar:TSE . The last column in Table 3 provides the mutation score (MS) for our case study subjects; it corresponds to the proportion of mutants detected by the test suite. The highest mutation score is observed with MLFS, whose test suite achieves MC/DC adequacy chilenski1994applicability . The lowest mutation score is observed with ASN1lib, which is automatically generated by ASN1SCC using a grammar-based approach ASN1:paper . Our subjects’ mutation scores are in line with empirical investigations reporting mutation scores ranging from 55% to 95% Ramler2017 ; delgado2018evaluation , for CPS software.
@@
The number of live mutants for each subject is: 443 for LIBU, 1,347 for ASN1lib, 3,891 for MLFS, 581 for ESAIL, and 99 for S5. For LIBU, MLFS, and ESAIL, we configured MOTIF to generate arrays of a specific size and, for void pointers (i.e., void *), to create variables or arrays considering the expected data type. Regarding S5, the relatively limited number of mutants considered is due to the need for the manual editing of most of the fuzz drivers generated by MOTIF (see Section 4). Indeed, advanced C++ features, including abstract classes, templates, friend functions, and unnamed namespaces, as well as the standard library implementations, render the drivers automatically generated by MOTIF ineffective. For example, vector<>, one of the standard template libraries (STL), uses pointers to trace the stored data, which renders the memcopy-based solution adopted by MOTIF to instantiate objects inappropriate (pointer addresses cannot be filled with data generated by the fuzzer). To address such limitation, we manually modified the fuzzing drivers automatically generated by MOTIF. Precisely, in total, to kill 99 mutants, MOTIF generated 25 drivers, 21 of which had to be manually modified. However, 14 drivers required a simple modification such as instantiating a template class (e.g., vector<int>), while seven drivers required an initialization step for the class under test, which we copied from the unit test cases provided by S5 developers. In addition, recall that, as discussed in Section 4.1, the need for such manual modifications is a limitation of the MOTIF implementation, not the approach, and can be overcome through an improved static analysis step. Further, compared to manual mutation testing, the automated generation of both the fuzzing driver and the test utility class still significantly decreases manual effort.

### v2

@@ 1 Introduction
We propose MutatiOn TestIng with Fuzzing (MOTIF), an approach that automatically generates fuzz drivers, leverage the data produced by fuzzers, and exercise a mutated function and the corresponding original function, looking for diverging outputs to detect mutants. MOTIF automatically generates seed files for fuzzers and integrates strategies to process mutant-killing inputs to eliminate false positives due to nondeterminism; in this context, possible sources of nondeterminism are functions interacting with timers, accessing the file system, or receiving inputs from sensors.
Instead of designing a dedicated fuzzing algorithm, MOTIF leverages state-of-the-art (SOTA) fuzzers, as their bucket-based coverage strategy has demonstrated being effective in maximizing the exploration of different execution paths. This strategy not only selects inputs that lead to different software states, which could help killing mutants, but enables tracking the behavioral differences between the original and mutated functions since such differences lead to distinct execution paths that can be exercised only by mutant-killing inputs. Although MOTIF can target any C/C++ program, we designed it taking into account characteristics that are prominent in embedded software for CPS. Specifically, CPS software is not executed from the textual console but directly within the operating system, largely relies on floating point instructions, and leverages specific OS features (e.g., RTEMS) and libraries that complicate its analysis or symbolic interpretation. For these reasons, we present MOTIF as a tool for C/C++ embedded software for CPS.
@@
  • Performing a large empirical study to investigate (RQ1) the best fuzzing configurations (compiler, sanitizers, and coverage metric) for mutation testing. Note that the identification of the best fuzzing configuration for different application contexts remains an open problem EURECOM7801 . We thus focused on identifying the best configurations for safety-critical space software implemented in C and C++, including mathematical libraries, control software, and data processing software; our results shall thus generalize to other types of CPS including these kinds of components, for example, avionic and robotic CPS. Our experiments show that the best results are obtained when relying (1) on the Clang compiler, which leads to the quickest fuzz drivers, (2) the address sanitizer (ASAN), which prevents false positives due to violation of function preconditions, and (3) the LAF-Intel LAF:Blog optimization for coverage, which enables computing a better fitness score for mathematical functions, which are, in turn, among the most critical software components. Specifically, the best fuzzing configuration, compared to a standard fuzzing configuration (i.e., AFL++ aflchurn with the GCC GCC compiler), leads to an improvement varying between 1.53 and 4.34 percentage points. Although limited, such improvements may still enable the detection of critical faults; for example, when reusing the generated test cases for regression testing. Further, our findings help engineers make informed decisions on projects setups. For example, mutation testing improvements may not justify a large development endeavor to enable a project to be recompiled with Clang. However, when compiling projects with Clang, ASAN, and LAF-Intel is feasible (e.g., it is sufficient to change a few configuration parameters), the best configuration is a better choice (e.g., enables testing additional boundary cases) and might be considered also for other fuzzing applications beyond mutation testing.
  • Assessing (RQ4) the performance gain achieved by relying on hybrid-fuzzing, which combines fuzzing and symbolic execution. Our results confirm the findings of related work that demonstrate that hybrid fuzzing leads to improved test effectiveness; however, in our experiments, its applicability is limited as it helps significantly improve the mutation score for only one in four subjects. The key advantage observed with hybrid-fuzzing is the reduction of the median time required to kill a mutant in three out of four subjects (between 14% and 32% faster), which suggests using hybrid-fuzzing for large projects leading to a large number of mutants.
  • Evaluating (RQ5) the cost-effectiveness of reusing mutant-killing inputs to test live mutants; our results suggest that reusing mutant-killing inputs to kill other mutants is recommended when there is a sufficient test budget of at least 5000 seconds per mutant in our subjects and, possibly, in similar systems.
@@ 3.4 Step 4 – Perform mutation testing
In Step 4, MOTIF runs the fuzzer to generate inputs for the executable fuzzing driver until either the time budget expires or a test input that can kill the mutant is found; in our experiments we leverage the AFL++ fuzzer but the approach is generic and can work with any grey-box fuzzer for C/C++. The fuzzer keeps generating input files until it reports one or more crashes, after which MOTIF halts the fuzzer. This process leads to the generation of fuzzing driver logs and crashing inputs (i.e., input files that caused a crash during the execution of the fuzzing driver). Since fuzzers generate several input files from each input taken from the file pool, and since all of them are executed by the fuzzer, more than one crashing input may be reported.
@@
MOTIF retains all test inputs that kill a mutant. However, the function under test may produce non-deterministic outputs meaning that, despite observed differences in outputs, the inputs may not have actually killed the mutant. For instance, two consecutive invocations of a function that reads and writes global variables may lead to different outputs even if the mutated statement is not exercised; consequently, the input suggested by the fuzzer would be a false positive. To minimize false positives, MOTIF automatically re-executes every test input killing a mutant using a modified version of the fuzzing driver that invokes the original function instead of the mutated function (see the post-processing of Step 4 in Figure 1). If this false positive driver, as we refer to it, reports a difference in outputs of the two function calls, it implies that the function under test is non-deterministic and thus that the input may not kill the mutant. MOTIF considers mutants exclusively killed by false positive inputs to be live. To kill mutants in functions that modify global state variables, the user must manually add, in the fuzzing driver between the two function calls, instructions to reset each state variable, similar to other fuzzing approaches for unit and library testing (e.g., LibFuzzer LibFuzzer ).
@@ 3.5 Step 5 – Test the SUT
1// expected values
2const unsigned char ex_pVal[8056] = {0x03,0x00,0x00,0x00,0x01,0x00,0x00,...};
3const unsigned char ex_pErrCode[4] = {0x9B,0x01,0x00,0x00};
4const unsigned char ex_return[1] = {0x00};
5
6int main(int argc, char** argv){
@@
15 // assert statements
16 assert(0==compare_value((char *)&pVal, ex_pVal, sizeof(pVal)));
17 assert(0==compare_value((char *)&pErrCode, ex_pErrCode, sizeof(pErrCode)));
18 assert(0==compare_value((char *)&_return, ex_return, sizeof(_return)));
[…]
@@
In this Step, MOTIF generates a test driver for the SUT. An example test driver for function T_POS_IsConstraintValid is shown in Listing 3. The test driver is identical to the fuzzing driver except that (1) it only invokes the original function (Line 10 in Listing 3), (2) instead of comparing outputs from two function invocations, it prints the output data generated by the original function (Lines 12 to 14), and (3) it includes assertions that compare the execution results of the original function with the expected results (Lines 15 to 18). The expected values used in the assertion statements are those observed during fuzzing for the original function and are stored in the test driver as a char array with a variable whose name starts with ex_ (Lines 2 to 4).
The test driver is used to test the original SUT with the inputs that kill the mutant. To proceed with testing, engineers shall verify if the generated assertions are correct, which is a typical activity for automatically generated test cases. However, since our implementation of MOTIF stores expected values in binary form, we find that the easiest way to verify assertion correctness is by executing the test case and observing its outputs. Precisely, engineers shall observe the values printed out by the test driver for each variable of interest (Lines 12 to 14 in Listing 3); note that, sometimes, with complex data structures, the inspection of values is simplified through the use of a debugger. If the test outputs match the specifications, then the software behaves correctly; otherwise, the test case has discovered a fault. This scenario highlights one of the key advantages of mutation testing: When testing the SUT with inputs that detect simulated human mistakes (mutants), actual faults in the SUT are more likely to be discovered than with randomly selected inputs (e.g., in our experiments we discovered five faults in our subjects). When a fault is discovered, engineers need to fix the software but also update the assertions to reflect the expected behaviour instead of the behaviour observed during fuzzing. Since the assertions generated by MOTIF match the output observed during fuzzing, the test case is expected to pass even if the output does not match the specifications. If the test case does not pass, it means that the software behaves non-deterministically and MOTIF did not discover such nondeterminism automatically (it never happened in our experiments); in such case, engineers shall decide whether to keep the test case and update the assertions manually or simply discard the test case (e.g., because it does not kill the mutant). If the test case passes and the engineers confirm that the software outputs are correct, it can be integrated into the software test suite for future regression testing.
@@ 4 MOTIF C++ extensions
15 return getCharge()/getAvgConsumption();
@@
27 double getAvgConsumption(){
28 return super::getAvgConsumption()-getAvgCharging();
@@ 32 ChargeableModule(double consumptionBoundary, double chargeBoundary){
33 this->consumptionBoundary = consumptionBoundary;
34 this->chargeBoundary = chargeBoundary;
@@ 55 StaticMission(double presetDuration){
64template <class ObjectiveT>
65class DynamicMission: Mission {
@@
Inheritance enables a class to reuse (i.e., inherit) the methods declared by its superclass, if any. Inheritance complicates the unit testing automated by MOTIF-C because it leads to dependencies between superclasses and subclasses. Indeed, not only a subclass may execute superclass methods but the execution of a method declared by a superclass may trigger the execution of a method in a subclass. For example, the execution of method ChargeableModule.isAutonomous() in Listing 4 triggers the execution of the superclass method SatelliteModule.getAutonomy(), which, in turn, triggers the execution of ChargeableModule.getAvgConsumption() and ChargeableModule.getAvgCharging(). As a consequence, the solution adopted by MOTIF-C to test mutants (i.e., creating a copy of the original method and renaming it) would prevent the testing of methods following such pattern; for example, renaming a mutant of method getAvgConsumption() in ChargeableModule as mut_getAvgConsumption() would prevent its call from method SatelliteModule.getAutonomy(), thus rendering testing ineffective.
@@
5 double ctor_1;
6 double ctor_2;
@@ 13 TypeA* object;
17 get_value(&ctor_1, sizeof(ctor_1) );
18 get_value(&ctor_2, sizeof(ctor_2) );
@@
21 param_1 = new StaticMission(manual_1);
22
23 //instantiate the class under test
24 object = new TypeA(ctor_1, ctor_2);
@@
26 _return = object->isAutonomous(param_1);
@@
32 ret += compare_value(&this->ctor_1, &rhs->ctor_1);
33 ret += compare_value(&this->ctor_1, &rhs->ctor_2);
34 ret += compare_value(&this->param_1, &rhs->param_1);
35 ret += compare_value(&this->_return, &rhs->_return);
@@
In MOTIF::TestUtil, methods call and compare take care of the execution of the method under test and the comparison of the results obtained by the original and the mutated method, respectively. Method call (Lines 15-27, Listing 8), which executes the method under test, includes automatically generated instructions to populate parameters with fuzzed data (Lines 17 and 18). If the constructor of a function under test receives as input an abstract type, the instructions needed to construct the concrete type to be used shall be manually added (Lines 19-21). Last, method call invokes the method under test after instantiating the class under test and stores the returned value (Lines 23-26). Method compare (Line 29-37) compares the results obtained by the original and the mutated class. The fuzzing driver instantiates two utility classes for each original and mutated method (Lines 6 and 10, Listing 7), and invokes the call method for them (Lines 7 and 11, Listing 7). These results are compared by invoking the compare method (Line 13, Listing 7). Its logic follows the one of the fuzzing driver generated by MOTIF-C. Approaches for Java MU2 ; fraser2011mutation can leverage programming conventions and reflection to compare object states by leveraging the equals method, which is normally overridden to compare two instances of a class, or inspector methods, which can be identified through naming conventions or purity analysis (simpler to achieve on Java bytecode Salcianu2005 ).
To deal with protected and private methods (i.e., encapsulation), instead of modifying the class under test to specify the test utility class as a friend class, we suggest the end-user to exercise, in the fuzzing driver, a public method that invokes the private/protected method under test. For example, to test mutant 2 in Listing 6, the end-user may still rely on the fuzzing driver shown in Listing 7, because it tests method isAutonomous, which, in turn, executes method getAvgConsumption (i.e., the mutated method). In Java, protected and private methods can be invoked through reflection Arcuri2017 .
@@ 4.1 Manual tuning of MOTIF’s test utility class
Like MOTIF-C, MOTIF-C++ automatically identifies the variables to create in the test utility class based on method signatures. However, MOTIF-C++ requires the manual specification of the class to instantiate to deal with parametric and subtype polymorphism. For example, if a method declares parameters of an abstract type (e.g., isAutonomous receives one parameter of the abstract type Mission), the end-user needs to specify the additional parameters required to instantiate the concrete class to be used to exercise the method under test (Line 9, Listing 8).
@@ 5.1 Subjects of the study
Our software subjects are provided with test suites whose code coverage is reported in Table 3. Most test suites do not achieve 100% statement coverage because they include components that are tested with specific hardware not available to us; therefore, we generated mutants only for the covered statements. We generated mutants with MASS Oscar:TSE ; MASSTOOL ; specifically, we rely on all the mutation operators supported by MASS (i.e., the sufficient set offutt1996experimental and the deletion set delamaro2014designing ; andrews2005mutation ), which proved effective in previous experiments on similar subjects. The mutation operators implemented by MASS are shown in Table 4. Though MASS does not include mutation operators specific to object-oriented programming languages delgado2017assessment , it is the only tool that enabled scalable mutation analysis in all our subjects, based on our previous experiments Oscar:TSE ; further, including C++-specific operators would have led to a different setup across subjects. In addition, MASS automatically excludes mutants that are identified as equivalent or duplicate according to trivial compiler equivalence methods [^4] The percentage of equivalent and duplicate mutants excluded by MASS before computing the MS is significant: 23.93% for MLFS, 28.00% for LIBU, 26.19% for ASN1lib, 36.10% for ESAIL, 23.29% for S5. Oscar:TSE . The last column in Table 3 provides the mutation score (MS) for our case study subjects; it corresponds to the proportion of mutants detected by the test suite. The highest mutation score is observed with MLFS, whose test suite achieves MC/DC adequacy chilenski1994applicability . The lowest mutation score is observed with ASN1lib, which is automatically generated by ASN1SCC using a grammar-based approach ASN1:paper . Our subjects’ mutation scores are in line with empirical investigations reporting mutation scores ranging from 55% to 95% Ramler2017 ; delgado2018evaluation , for CPS software.
Table 4: Mutation operators implemented in MASS.
Operator Description
Sufficient Set
ABS $\{(v,-v)\}$
[…]
ICR $\{i,x)\,|\,x\in\{1,-1,0,i+1,i-1,-i\}\}$
[…]
SDL $\{(s,\texttt{remove}(s))\}$
UOI $\{(v,\texttt{--}v),(v,v\texttt{--}),(v,\texttt{++}v),(v,v\texttt{++})\}$+ ) }
OODL
AOD $\{((t_{1}\,op\,t_{2}),t_{1}),((t_{1}\,op\,t_{2}),t_{2})\,|\,op\in\{\texttt{+},\texttt{-},\texttt{*},\texttt{/},\texttt{\%}\}\}$
LOD $\{((t_{1}\,op\,t_{2}),t_{1}),((t_{1}\,op\,t_{2}),t_{2})\,|\,op\in\{\texttt{\&\&},||\}\}$& , | | } }
[…]
BOD $\{((t_{1}\,op\,t_{2}),t_{1}),((t_{1}\,op\,t_{2}),t_{2})\,|\,op\in\{\texttt{\&},\texttt{|},\land\}\}$
SOD $\{((t_{1}\,op\,t_{2}),t_{1}),((t_{1}\,op\,t_{2}),t_{2})\,|\,op\in\{\texttt{>>},\texttt{<<}\}\}$> , << } }
Other
[…]
@@
Table 5: The number of live mutants for each subject and their distribution for each mutation operator obtained from MASS.
Operators LIBU ASN1lib MLFS ESAIL S5 Total
ABS 6 (01.37%) 20 (01.48%) 232 (05.96%) 0 (00.00%) 5 (05.05%) 263 (04.14%)
AOR 7 (01.60%) 0 (00.00%) 369 (09.48%) 201 (34.60%) 14 (14.14%) 591 (09.30%)
ICR 56 (12.81%) 426 (31.63%) 809 (20.79%) 103 (17.73%) 39 (39.39%) 1433 (22.55%)
LCR 52 (11.90%) 18 (01.34%) 96 (02.47%) 5 (00.86%) 0 (00.00%) 171 (02.69%)
ROR 115 (26.32%) 303 (22.49%) 348 (08.94%) 48 (08.26%) 7 (07.07%) 821 (12.92%)
SDL 33 (07.55%) 197 (14.63%) 196 (05.04%) 117 (20.14%) 11 (11.11%) 554 (08.72%)
UOI 108 (24.71%) 186 (13.81%) 1377 (35.39%) 83 (14.29%) 10 (10.10%) 1764 (27.76%)
AOD 0 (00.00%) 0 (00.00%) 234 (06.01%) 11 (01.89%) 6 (06.06%) 251 (03.95%)
LOD 40 (09.15%) 39 (02.90%) 5 (00.13%) 3 (00.52%) 0 (00.00%) 87 (01.37%)
ROD 19 (04.35%) 126 (09.35%) 72 (01.85%) 6 (01.03%) 0 (00.00%) 223 (03.51%)
BOD 1 (00.23%) 0 (00.00%) 48 (01.23%) 2 (00.34%) 0 (00.00%) 51 (00.80%)
SOD 0 (00.00%) 0 (00.00%) 11 (00.28%) 0 (00.00%) 0 (00.00%) 11 (00.17%)
LVR 0 (00.00%) 32 (02.38%) 94 (02.42%) 2 (00.34%) 7 (07.07%) 135 (02.12%)
Total 437 1347 3891 581 99 6355
Table 5 reports the number of live mutants for each subject and the distribution of mutants across different mutation operators, as generated by MASS. The total number of mutants per subject is 437 for LIBU, 1,347 for ASN1lib, 3,891 for MLFS, 581 for ESAIL, and 99 for S5. While these mutants show that all types of mutation operators are covered, certain operators, such as ICR and UOI, contribute a larger number of mutants than others. This imbalance arises from both the nature of the mutation operators and the characteristics of the subjects. In particular, when test suites inadequately cover boundary conditions, mutants generated through small increments or decrements (e.g., by adding or subtracting 1) are more likely to survive. For example, in our case, many mutants deal with arrays. Modifying array indices by +1 or -1 often does not significantly affect program output, allowing such mutants not to be detected by the test suite.
For LIBU, MLFS, and ESAIL, we configured MOTIF to generate arrays of a specific size and, for void pointers (i.e., void *), to create variables or arrays considering the expected data type. Regarding S5, the relatively limited number of mutants considered is due to the need for the manual editing of most of the fuzz drivers generated by MOTIF (see Section 4). Indeed, advanced C++ features, including abstract classes, templates, friend functions, and unnamed namespaces, as well as the standard library implementations, render the drivers automatically generated by MOTIF ineffective. For example, vector<>, one of the standard template libraries (STL), uses pointers to trace the stored data, which renders the memcopy-based solution adopted by MOTIF to instantiate objects inappropriate (pointer addresses cannot be filled with data generated by the fuzzer). To address such limitation, we manually modified the fuzzing drivers automatically generated by MOTIF. Precisely, in total, to kill 99 mutants, MOTIF generated 25 drivers, 21 of which had to be manually modified. However, 14 drivers required a simple modification such as instantiating a template class (e.g., vector<int>), while seven drivers required an initialization step for the class under test, which we copied from the unit test cases provided by S5 developers. In addition, recall that, as discussed in Section 4.1, the need for such manual modifications is a limitation of the MOTIF implementation, not the approach, and can be overcome through an improved static analysis step. Further, compared to manual mutation testing, the automated generation of both the fuzzing driver and the test utility class still significantly decreases manual effort.